OpenStudy (basketball12345):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, -7)???

OpenStudy (dumbcow):

\[y = ax^2\] \[a = \frac{1}{4p}\] where "p" is distance from focus to vertex distance between (0,0) and (0,-7) is (-7 - 0) = -7 therefore p = -7

OpenStudy (basketball12345):

-7x? is one of the answers

OpenStudy (dumbcow):

No plug -7 in for p to get the correct "a" coefficient